OpenStudy (anonymous):

simplify this. 1/(x-1)+1/(4x) = 3 can you show me how you did it?

OpenStudy (anonymous):

you multiply through right?

OpenStudy (angela210793):

u put them in the same denom which is (4x)*(x-1) (4x-x+1)/4x*(x-1)=3

OpenStudy (anonymous):

how did you get the first part?

OpenStudy (angela210793):

by multiplying 1/(x-1) with 4x up and down and by multiplying by(x-1) up and down (1/4x) did u get it?

OpenStudy (anonymous):

sort of.

OpenStudy (anonymous):

so you get 4x+x-1 = 3?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

4x/4x-4+x-1/4x^2-4=3?

OpenStudy (angela210793):

all u have to do is multiply wht u have in denom and then from wht u get in the left crossmultiply

OpenStudy (anonymous):

so confusing.

OpenStudy (angela210793):

no it is not confusing.....just focus....multiply 4x(x-1) in denom and u get 4x^2-4x (3x+1)/(4x^2-4x)=3 now crossmultiply 3x+1=3(4x^2-4x)

OpenStudy (anonymous):

are you multiplying the top and the bottom by (4x)(x-1)?

OpenStudy (angela210793):

noo see here a/b+c/d=(ad+bc)/bd understand it?

OpenStudy (anonymous):

lol I guess I understand the formula..

OpenStudy (anonymous):

in this case it would be (1+4x)+(x-1+1)/(x-1)(4x)

OpenStudy (angela210793):

in this case would be[ 1*4x-(x-1)*1]/(x-1)(4x)

OpenStudy (anonymous):

3x-4/(x-1)(4x)? lol

OpenStudy (anonymous):

I'm sorry. I'm just going to skip this question for now.

OpenStudy (anonymous):

thanks for trying though.

OpenStudy (angela210793):

I'm sorry i wish i could explain u better,how abt twiddla?